OpenStudy (anonymous):

what is the best estimate for this product? 81 x409

OpenStudy (anonymous):

81*409=33,129

OpenStudy (anonymous):

For estimation do 33,130

OpenStudy (solomonzelman):

80 × 400 = 8 × 4 and then write 3 zeros in front of it.

OpenStudy (anonymous):

Or 33,000

OpenStudy (solomonzelman):

No, estimating the numbers you multiply before multiplying. This is silly, but that is how it works .

OpenStudy (solomonzelman):

Like I did, I estimated 81≈80 409≈400

OpenStudy (anonymous):

Which way is correct SolomonZelman?

OpenStudy (solomonzelman):

Well, I would not round up the numbers before multiplying, I agree that it is stupid, but that is what they are asking for here -:(

OpenStudy (solomonzelman):

They are teaching a common sense here, that if you would try to multiply 28 and 401, you would say it is about 12,000. Because you would multiply 30 × 400. Even though the exact answer would be 11,228. This is done, so that people can find what the number can be when they have no calculator or paper.

OpenStudy (anonymous):

Thank you! You taught me something really valuable.

OpenStudy (solomonzelman):

Well, I personally would multiply those together or use a calculator on an i-phone. People sometimes use this method, but I am not a fan of it. What YOU did is better, estimating after multiplying, not before.